40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Jun 2 10:28:26 2022, Yehonathan, Update, WOPO, Installing 1811 PDs 
    Reply  Wed Jul 20 15:55:41 2022, Yehonathan, Update, WOPO, Installing 1811 PDs Diabolo_Current_Power_at_PDs.png
       Reply  Tue Aug 23 22:23:06 2022, awade, Update, WOPO, Installing 1811 PDs 
Message ID: 2796     Entry time: Wed Jul 20 15:55:41 2022     In reply to: 2777     Reply to this: 2800
Author: Yehonathan 
Type: Update 
Category: WOPO 
Subject: Installing 1811 PDs 

{Yehonathan, Paco}

We comfirmed that the DC ouput of one of the 1811s is bad. We set out to measure the AC response of the PDs.

For this, we decided to use the current modulation on the Diabolo laser which is rated to have 0.1 A/V and a bandwidth of 5kHz. We calibrated the current to optical power by swiping the current and measuring the power at the homodyne PDs using a power meter. The laser power before the 1064nm PZT mirror was measured to be 5mW.

Attachment shows the measurement and a linear fit with slope=0.97 mW/A.

We drove the current modulator using a sine wave from a function generator with 1kHz 0.5Vpp. When we looked on the PD AC signal port in the Oscilloscope we saw 2 Vpp 12Mhz signal. We passed the signal with a low pass filter but again we saw mostly noise.

We took the PDs to the 40m PD test stand but we accidently fried Jenne's laser.

Next, we should just use the Moku network analyzer instead of the scope to measure the response in the QIL using the Diabolo current modulator.

Attachment 1: Diabolo_Current_Power_at_PDs.png  21 kB  Uploaded Wed Jul 20 17:05:48 2022  | Hide | Hide all
Diabolo_Current_Power_at_PDs.png
ELOG V3.1.3-